OpenStudy (anonymous):

A monkey of mass 20kg is holding a vertical rope . the rope can break when a mass of 25kg is suspended from it.What is the maximum acceleration with which the monkey can climb up along the rope? Take g=10 meter per second square.

OpenStudy (anonymous):

Is it simple F=ma, where m=20kg, F=(25-20)*g=50(N)? So, a = 50/20 = 2,5 (m/s), I think so.

OpenStudy (anonymous):

yes thanxx